OpenStudy (anonymous):

Write an equation in slope-intercept form for a line that passes through the given pair of points. (4,-7), (-2,-10)

OpenStudy (anonymous):

x^1-y^1 x^2-y^2

OpenStudy (anonymous):

lol i dont get it

OpenStudy (anonymous):

What about this one Write an equation in slope-intercept form for a line that passes through the given pair of points. (-6,5), (-9,0)

OpenStudy (anonymous):

@asib1214 help me lmao

OpenStudy (anonymous):

ohhhhh now i get it, thats the formula

OpenStudy (texaschic101):

first you have to find the slope by using the slope formula... slope(m) = (y2 - y1) / (x2 - x1) (4,-7)....x1 = 4 and y1 = -7 (-2,-10)..x2 = -2 and y2 = -10 now we sub slope(m) = (-10 - (-7) / (-2 - 4) slope(m) = (-10 + 7) / -6 slope(m) = -3/-6 slope(m) = 1/2 Now use y = mx + b slope(m) = 1/2 use either of your points... (4,-7)....x = 4 and y = -7 now we sub -7 = 1/2(4) + b -7 = 2 + b -7 - 2 = b -9 = b so your equation is : y = 1/2x - 9 However, we need it in standard form Ax + By = C y = 1/2x - 9 --- multiply everything by 2 2y = x - 18 -- subtract x from both sides -x + 2y = -18 -- multiply by -1 to make x positive x - 2y = 18 <== standard form

OpenStudy (anonymous):

lol give me a minute to absorb that

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (0 - 5) / (-9 - (-6) slope(m) = -5/(-9 + 6) slope(m) = -5/-3 slope(m) = 5/3

OpenStudy (texaschic101):

oops...you needed it in slope intercept form so the 1st one is : y = 1/2x - 9

OpenStudy (texaschic101):

can you do the 2nd one....we already figured the sloep to be 5/3

OpenStudy (anonymous):

i actually already knew how to do it, i just needed refresher. But i understand it perfectly now. Thank you very much sir

OpenStudy (texaschic101):

lol....not a sir...I am a lady

OpenStudy (anonymous):

i knew i should have just left it at tyvm and not asume anything...

OpenStudy (anonymous):

anyway, tyvm lady/gal/girl/madam which ever you prefer

OpenStudy (texaschic101):

lol...your very welcome sir :)

OpenStudy (anonymous):

One more! I have absolutely no idea how to do this Solve the equation or formula for the variable specified. .

OpenStudy (anonymous):

just realzed that the problem isnt even there...

OpenStudy (anonymous):

B = 7a^2b, solve for b